OpenStudy (mrs.ambrose614):

Question 20 (Multiple Choice Worth 1 points) (03.06 MC) A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $155 for 100 minutes of usage and $245 for 200 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x? Draw a graph which joins the points (155, 100) and (245, 200) and has a slope = 0.9 Draw a graph which joins the points (155, 100) and (245, 200) and has a slo

OpenStudy (michaelpersaus48):

b

OpenStudy (mrs.ambrose614):

Did you see the other awnsers???? Question 20 (Multiple Choice Worth 1 points) (03.06 MC) A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $155 for 100 minutes of usage and $245 for 200 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x? Draw a graph which joins the points (155, 100) and (245, 200) and has a slope = 0.9 Draw a graph which joins the points (155, 100) and (245, 200) and has a slope = 1.1 Draw a graph which joins the points (100, 155) and (200, 245) and has a slope = 1.1 Draw a graph which joins the points (100, 155) and (200, 245) and has a slope = 0.9

OpenStudy (michaelpersaus48):

i have answered this question before on peer answer so b

OpenStudy (will.h):

No direct answers if you know the answer then show us the way to it

OpenStudy (michaelpersaus48):

Draw a graph which joins the points (100, 135) and (500, 375) and has a slope = 0.6 is the one among the following that can best describe the steps to draw the graph of y against x. The correct option among all the options that are given in the question is the second option or option "B".

OpenStudy (mrs.ambrose614):

Is this correct will. Explains it clear

OpenStudy (will.h):

x represents minutues and y represnts the cost thus we have the points (100,155)(200,245) Slope = y2-y1/x2-x1 slope = 245 - 155/200 - 100 slope = 90/100 Slope = 0.9 so it would be D actually Draw a graph which joins the points (100, 155) and (200, 245) and has a slope = 0.9

OpenStudy (mrs.ambrose614):

The graphs of functions f(x) and g(x) = f(x) + k are shown below: graph of line f of x going through ordered pairs 0, 0 and 2, 4. Graph of line g of x going through ordered pairs 0, 2 and 1.5, 5. The value of k is ___. @MichaelPersaus48

OpenStudy (mrs.ambrose614):

The graphs of functions f(x) and g(x) = f(x) + k are shown below: graph of line f of x going through ordered pairs 0, 0 and 2, 4. Graph of line g of x going through ordered pairs 0, 2 and 1.5, 5. The value of k is ___. @ will.h

OpenStudy (will.h):

since k is about the y intercept we know that f(x) has a y intercept of 0 because that's when x = 0 and the y intercept of function g(x) is 2 therefor the change would be k 2 - 0 = 2 therefor k is 2

OpenStudy (mrs.ambrose614):

Would it be positive or negative 2????

OpenStudy (will.h):

positive